11 June 2014: Wednesday

LDAP:

  • The problem with former students that had dropped out of the Banner feed on April 30, were corrected yesterday and 15 former students were delivered.  The only one that had any issue was the one that we’d manually munged to work.
  • There was a single item match with a different last name.  Either the existing record or the new record has the first and last name reversed.  Checking.

BACKUPS:

  • CoM/IS clones finished.  The other two continue on.
  • The three tapes from yesterday (111302L4, 111303L4, 111304L4) were all found upon physical inspection to have their write protect tabs closed.  The tabs were opened and the tapes returned to the tape library.
  • Relabeling is complete on the 384 tapes that were returned from the vaults on Monday.

FootPrints:

  • Waiting for BMC to get back to me about installing 11.6.02 with Java 1.7.
  • BMC provides a workaround which is essentially an admission that Java 1.7 is fine and the update script is being overly cautious.

VMware backup rfp:

  • Processing, processing, processing

Tags:

Leave a Reply

You must be logged in to post a comment.